OpenStudy (anonymous):

The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work. The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work.

OpenStudy (anonymous):

This pretty much shows all of my work.

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

The image you posted is just a list of questions

OpenStudy (anonymous):

Yes, though I don't know how to do it...

OpenStudy (anonymous):

oh, i thought you made a mistake, have you tried them? where did u get stuck

OpenStudy (anonymous):

I frankly just don't understand this, I've tried reading through the lesson, and I don't seem to understand what I'm reading :(

OpenStudy (anonymous):

lets do the first one, so you have\[d = ut + 0.5a t^{2}\] Solving for a just means making a the subject of this formula I'll start you off, \[d = ut + 0.5a t^{2}\] tells us that \[d - ut = 0.5at^{2}\] so you see how I got that?

OpenStudy (anonymous):

While you are doing the first part of the question where they say "solve this formula for", try to forget about what each of the letter means and just treat it as a basic math problem

OpenStudy (anonymous):

Hm, so what I need to do is to turn the first operation into it's opposite, while the second operation becomes an equal sign? So, what is the reason they do that?

OpenStudy (anonymous):

@Crixpack